# BootOOD: Self-Supervised Out-of-Distribution Detection via Synthetic Sample Exposure under Neural Collapse

### v1

@@ 4.5 Ablations and Design Choices
We conduct ablations on CIFAR-100 (ResNet-18), averaging results over all near-/far-OOD test sets. We independently remove: (1) the Phase-1 warm-up, (2) the radius self-supervision losses (rad-cls & rad-reg), (3) the feature-separation loss, and (4) vary the number of radius bins $K$ used by the training-time radius head. Table 3 and Table 4 report relative changes against the full model.
@@
Removing the Phase-1 warm-up leads to a clear drop in AUROC (–2.6%) and a large increase in FPR@95% (+9.7%). This shows that letting the classifier first learn stable ID prototypes before turning on OOD regularization is important for both ID accuracy and OOD detection.
@@
Ablating the radius self-supervision ($\mathcal{L}_{\text{rad-cls}}+\lambda_{\text{mse}}\mathcal{L}_{\text{rad-reg}}$) hurts performance the most (–4.2% AUROC). Once these losses are removed, the model behaves close to plain CE training, and the feature norms no longer provide a strong ID vs. OOD separation signal.
@@
Dropping the feature separation term yields a smaller but consistent degradation (–0.7% AUROC, higher FPR@95%). Thus, separation is not the main driver but still gives a useful extra margin, especially for near-OOD close to class boundaries.
Effect of radius bins $K$.
We also vary the number of radius bins $K$ in the training-only radius head that supervises pseudo-OOD radii. As shown in Table 4, $K=1$ (pure regression without meaningful radius classes) performs worst, while increasing $K$ initially improves near-OOD AUROC and FPR@95 by providing a richer discretization of the inner-shell targets. Performance peaks at $K=4$ and then slightly degrades for $K=6$, suggesting that too many bins introduce unnecessary quantization noise without clear benefits. We therefore adopt $K=4$ as our default.
Table 4: Near-OOD ablation on the number of radius bins $K$ in the training-time radius head (CIFAR-100).
$K$ (radius bins) AUROC FPR@95%
$K=1$ 77.67($\pm 0.72$) 62.35($\pm 0.61$)
$K=3$ 80.46($\pm 0.39$) 56.47($\pm 0.41$)
$K=\textbf{4}$ 83.09($\pm 0.23$) 51.03($\pm 0.36$)
$K=6$ 81.88($\pm 0.16$) 53.58($\pm 0.26$)
@@
All components matter: Phase-1 warm-up provides stability, radius self-supervision is the main driver of OOD separability, feature separation adds complementary gains, and a moderate number of radius bins (K=4K{=}4) yields the best near-OOD performance.

### v2

@@ 4.5 Ablations and Design Choices
We conduct ablations on CIFAR-100 (ResNet-18), averaging results over all near-/far-OOD test sets. We study the impact of the following design choices: (1) the Phase-I Neural-Collapse warm-up, (2) the radius self-supervision mechanism, (3) the decomposition of radius self-supervision into classification-only and regression-only variants, (4) the feature-separation loss, (5) the number of radius levels $K$ used by the training-time radius head, and (6) the strength of OOD supervision controlled by $\lambda_{\text{ood\_max}}$. Tables 3 and 5 report component-removal ablations against the full model, while Tables 4 and 6 report sensitivity to the hyperparameters $K$ and $\lambda_{\text{ood\_max}}$.
@@
Removing the Phase-1 warm-up leads to a clear drop in AUROC and a large increase in FPR@95%. This shows that letting the classifier first learn stable ID prototypes before turning on OOD regularization is important for both ID accuracy and OOD detection.
@@
Ablating the radius self-supervision ($\lambda_{\text{cls}}\mathcal{L}_{\text{rad-cls}}+\lambda_{\text{mse}}\mathcal{L}_{\text{rad-reg}}$) hurts performance the most. Once these losses are removed, the model behaves close to plain CE training, and the feature norms no longer provide a strong ID vs. OOD separation signal.
@@
Dropping the feature separation term yields a smaller but consistent degradation. Thus, separation is not the main driver but still gives a useful extra margin, especially for near-OOD close to class boundaries.
Effect of radius levels $K$.
We vary the number of radius levels $K$ used by the training-only radius head, where each level corresponds to a distinct target radius (i.e., a concentric shell). As shown in Table 4, $K=1$ (degenerating to pure regression without meaningful radius classification) performs worst. Increasing $K$ initially improves near-OOD AUROC and FPR@95 by providing a richer set of discrete radial targets. Performance peaks at $K=4$ and then slightly degrades for $K=6$, suggesting that excessive radial levels introduce unnecessary complexity without clear benefits. We therefore adopt $K=4$ as the default setting for CIFAR-100 in the following experiments.
Table 4: Near-OOD ablation on the number of radius shells $K$ used by the training-time radius head (CIFAR-100).
[…]
$K=1$ -1.33 +5.14 -1.36
$K=3$ -0.54 +3.23 -0.65
$K=4$
$K=6$ -1.08 +1.25 -0.57
Ablation on CIFAR-100: joint vs. regression-only vs. classification-only.
Table 5 compares the full joint model with its single-branch variants. The classification-only variant learns discrete radial bins but lacks an absolute scale for $\lVert z-\mu\rVert$, leading to drift within bins and degraded AUROC, FPR@95, and ID accuracy. The regression-only variant enforces an absolute radius but misses the discretized margin and multi-scale gradients from the bins, further hurting detection. Combining both terms yields the best trade-off: bin-level separation from classification and radius refinement from regression, resulting in the strongest OOD separability and the smallest drops in ID-ACC.
Effect of $\lambda_{\text{ood\_max}}$ on CIFAR-100.
Table 6 reports the effect of varying $\lambda_{\text{ood\_max}}$ relative to the default value $0.1$. A smaller coefficient ($0.05$) weakens the radial and directional supervision from the pseudo-OOD branch, yielding slightly worse AUROC, higher FPR@95, and a noticeable drop in ID accuracy. Increasing $\lambda_{\text{ood\_max}}$ ($0.2$ or $0.5$) makes the auxiliary loss overly dominant, over-contracting ID features and perturbing classifier directions, further degrading both AUROC and ID-ACC while increasing FPR@95. Overall, $\lambda_{\text{ood\_max}}=0.1$ provides the best trade-off between enforcing OOD-specific supervision and preserving the geometry of ID decision regions.
@@
Overall, all components play complementary roles. The Phase-I warm-up is crucial for reaching a stable Neural-Collapse regime before applying OOD-specific supervision. Radius self-supervision is the primary driver of OOD separability, and the joint formulation that combines radial classification with regression consistently outperforms either branch alone. Feature separation provides additional gains by reducing spurious class alignment. A moderate number of radius levels (K=4K{=}4) yields the best near-OOD performance on CIFAR-100, while the supervision strength must be carefully balanced: $\lambda_{\text{ood\_max}}{=}0.1$ offers the best trade-off between enforcing OOD-aware geometry and preserving in-distribution decision structure.
Table 5: Ablation on CIFAR-100: relative change (%) w.r.t. the joint (full) model.
[…]
Joint (full)
Classification only -0.65 +1.84 -1.92
Regression only -0.88 +4.84 -0.87
Table 6: Effect of $\lambda_{\text{ood\_max}}$ on CIFAR-100 (relative change w.r.t. $\lambda_{\text{ood\_max}}=0.1$).
[…]
0.10
0.05 -0.32 +0.61 -0.77
0.20 -0.78 +2.76 -1.23
0.50 -1.26 +4.25 -1.41